# Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices

### v2

@@ 4.2.2 Intuitionistic partial (level) valuations and intuitionistic truth tables
[…]
@@
  • if $\neg\alpha\in\Lambda$ then $\tilde{v}_{p}(\neg\alpha)\in\neg^{\mathcal{A}_{IPL}}(\tilde{v}_{p}(\alpha))$;
[…]
@@
Let $\Lambda\in FCS(\Omega)$. A partial valuation $\tilde{v}_{p}\in iPV(\Lambda)$ is an intuitionistic partial level valuation in $\mathcal{M}_{IPL}$ over $\Lambda$ iff
  1. $\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$, there exists an intuitionistic partial level valuation $\tilde{w}_{p}$ in $\mathcal{M}_{IPL}$ over $\Lambda$ such that $\tilde{w}_{p}(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $\tilde{w}_{p}(\beta)=\mathbf{T}$ whenever $\tilde{v}_{p}(\beta)=bT$.
@@
Given $v,w$ and $\alpha$ let $\mathsf{P}_{\Lambda}^{i}(v,w,\alpha)$ iff $w(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $w(\beta)=\mathbf{T}$ whenever $v(\beta)=\mathbf{T}$. If $\Lambda\in FCS(\Omega)$ then
@@
$\begin{array}[]{lll}iPLV(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV(\Lambda)$}\big{)}\}.\end{array}$
As observed in Remark 3.7 for $\mathcal{M}_{S4}$, since $\Lambda$ is finite then $iPV(\Lambda)$ is finite, hence the definition of $iPLV(\Lambda)$ is not cyclic.
@@
An intuitionistic truth table in $\mathcal{M}_{IPL}$ is an exhaustive list of all intuitionistic partial level valuations for a given finite domain closed under subformulas.
@@
The following is an example of the algorithm for searching the (intuitionistic) partial level valuations. [^5] This example, as the ones presented in Appendix A, were automatically generated by a Coq module for RNmatrices written by R. Leme, see  https://github.com/renatoleme/Forest We start with the full table for $\neg\neg(p\lor\neg p)$.
Table 4: Initial table.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) F U F U U
(2) F U F U T
(3) F U F T F
(4) F T T F U
(5) F T T F T
(6) T F T F U
(7) T F T F T
Table 5: First cycle.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) $\times$ 6, 7 $\times$ 4, 5, 6, 7 3
(2) $\times$ 7 $\times$ 5, 7 $\times$
(3) $\times$ $\emptyset$ $\times$ $\times$ $\times$
(4) $\times$ $\times$ $\times$ $\times$ $\emptyset$
(5) $\times$ $\times$ $\times$ $\times$ $\times$
(6) $\times$ $\times$ $\times$ $\times$ $\emptyset$
(7) $\times$ $\times$ $\times$ $\times$ $\times$
The cells marked with the multiplication symbol ($\times$) in a collumn $\alpha$ of a partial valuation $v$ indicates that $v(\alpha)\neq\mathbf{U}$, so we can ignore that cell. The cells marked with the empty set symbol ($\emptyset$) in a collumn $\alpha$ of a partial valuation $v$ indicates that $v(\alpha)=\mathbf{U}$ and there is no other line $w$ to validate $v$, so that line must be removed.
Table 6: Intermediate table.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) F U F U U
(2) F U F U T
(5) F T T F T
(7) T F T F T
Table 7: Second cycle.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) $\times$ 7 $\times$ 5, 7 $\emptyset$
(2) $\times$ 7 $\times$ 5, 7 $\times$
(5) $\times$ $\times$ $\times$ $\times$ $\times$
(7) $\times$ $\times$ $\times$ $\times$ $\times$
Table 8: Final table.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(2) F U F U T
(5) F T T F T
(7) T F T F T
Table 9: Validators.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(2) $\times$ 7 $\times$ 5, 7 $\times$
(5) $\times$ $\times$ $\times$ $\times$ $\times$
(7) $\times$ $\times$ $\times$ $\times$ $\times$
Note that, in the validators table, there is no occurrence of $\emptyset$. So, there is no further row to remove and we can stop the searching algorithm.
4.2.3 Soundness and completeness of the intuitionistic truth tables
@@
Now, it will be shown that the generalized notion of truth tables for IPL proposed in Subsection 4.2.2 is sound and complete. As in the case of S4, this will be shown by proving first analyticity and co-analyticity w.r.t. level valuations in $\mathcal{M}_{IPL}$. In order to do this it is considered, adapting [4, Definition 3.12], the intermediary notion of intuitionistic partial’ level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$.
@@
Let $\Lambda\in CS(\Omega)$. An intuitionistic partial valuation $\tilde{v}_{p}\in iPV(\Lambda)$ is an intuitionistic partial’ level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$ iff
@@
Recall the predicate $\mathsf{P}_{\Lambda}^{i}(v,w,\alpha)$ introduced in Remark 4.27. Clearly,
$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}_{IPL}$}\big{)}\}.\end{array}$
@@
It is an adaptation of the proof of [4, Lemma 3.13]. Suppose, by contradiction, that $\tilde{v}_{p}\notin iPLV^{\prime}(\Lambda)$. Let $\Gamma:=\{\beta\in\Lambda\mid\tilde{v}_{p}(\beta)=\mathbf{T}\}$. Then, there exists $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$ and, for every $w\in\mathcal{L}_{IPL}$, $w(\alpha)=\mathbf{F}$ implies that $w(\beta)\neq\mathbf{T}$ for some $\beta\in\Gamma$. By contraposition, this means that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$. But $v(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$ and $v(\alpha)=\mathbf{U}$, a contradiction. From this we infer that $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$. ∎
@@
Observe that, as it was done in Remark 3.22 for S4, it is possible to define an enumeration $\alpha_{1},\alpha_{2},\ldots\alpha_{m}\ldots$ (for $m\in\omega^{2}$) of $For(\Omega)$ such that $\textsf{c}(\alpha_{i})\leq\textsf{c}(\alpha_{j})$ if $i\leq j$ and, for every $i$ such that $\textsf{c}(\alpha_{i})>0$, if $\beta$ is a strict subformula of $\alpha_{i}$ then $\beta=\alpha_{j}$ for some $j<i$. Such an enumeration of $For(\Omega)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 00 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
@@
It is an adaptation of the proof of [4, Lemma 3.14] to $\mathcal{R}(\mathcal{M}_{IPL})$.

### v4

@@ 9.1 Intuitionistic partial (level) valuations and intuitionistic truth-tables
$CS(\Omega)=\{\Lambda\subseteq For(\Omega)\mid\mbox{ $\Lambda$ is non-empty and closed under subformulas}\},$
@@
  • if $\neg\alpha\in\Lambda$ then $\tilde{v}_{p}(\neg\alpha)\in\neg^{IPL}(\tilde{v}_{p}(\alpha))$;
  • if $#\in\{\to,\vee,\land\}$ and $\alpha\#\beta\in\Lambda$ then $\tilde{v}_{p}(\alpha\#\beta)\in\#^{IPL}(\tilde{v}_{p}(\alpha),\tilde{v}_{p}(\beta))$.
@@
Let $\Lambda\in FCS(\Omega)$. An intuitionistic partial level valuation in $\mathcal{M}_{IPL}$ over $\Lambda$ is a partial valuation $\tilde{v}_{p}\in iPV(\Lambda)$ satisfying the following:
  1. $\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$, there exists an intuitionistic partial level valuation $\tilde{w}_{p}$ in $\mathcal{M}_{IPL}$ over $\Lambda$ such that $\tilde{w}_{p}(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $\tilde{w}_{p}(\beta)=\mathbf{T}$ whenever $\tilde{v}_{p}(\beta)=\mathbf{T}$.
@@
Given $v,w$ and $\alpha$ let $\mathsf{P}_{\Lambda}(v,w,\alpha)$ iff $w(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $w(\beta)=\mathbf{T}$ whenever $v(\beta)=\mathbf{T}$. If $\Lambda\in FCS(\Omega)$ then
@@
$\begin{array}[]{lll}iPLV(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV(\Lambda)$}\big)\}.\end{array}$
Remark 9.5.
As in the case of S4 discussed in Remark 5.6, the impredicative definition of $iPLV(\Lambda)$ does not lead to circularity. Indeed, by using an algorithm (see Algorithm 1 below) similar to the Algorithm 3.4 given in [12, p. 15] for $\mathcal{M}_{S4}$, Definition 9.3 can be transformed into a predicative one (see Definition 9.11 and Proposition 9.19). This is the consequence of the fact that $iPLV(\Lambda)$ is obtained as the output of Algorithm 1, being the unique subset of $iPV(\Lambda)$ satisfying the condition of Remark 9.4, as it will be shown in Proposition 9.7. The set $iPLV(\Lambda)$ is precisely a (intuitionistic) truth-table, in the wider sense proposed here, based on $\mathcal{R}(\mathcal{M}_{IPL})$. It is obtained by means of a mechanical procedure, where its elements correspond to the rows of the table, and its columns correspond to the values assigned in each row to the elements of $\Lambda$.
Notation 9.6.
Given $\Lambda\in FCS(\Omega)$ with $n\geq 1$ elements, consider an arrangement $\alpha_{1},\ldots,\alpha_{n}$ of this set such that the complexity of $\alpha_{i}$ is less than or equal to the complexity of $\alpha_{j}$ if $i<j$ (so the atomic formulas occurring in $\Lambda$ appear as the first elements of the sequence, in arbitrary order). In this way, any subformula of $\alpha_{j}$ appears as $\alpha_{i}$ for some unique $i\leq j$. Suppose that the set $iPV(\Lambda)$ has $m\geq 1$ elements. Consider now an arrangement $v_{p}^{1},\ldots,v_{p}^{m}$ of its elements, ordered by lexicographical order, by stipulating that $\mathbf{F}<\mathbf{U}<\mathbf{T}$. In this way, $iPV(\Lambda)$ can be represented as a 2-dimensional array (i.e., an array of arrays, or a matrix) $V$ such that each row $V[i]$, for $1\leq i\leq m$, corresponds to the ithith element $v_{p}^{i}$ of $iPV(\Lambda)$. In turn, $V[i][j]$ is the truth-value assigned to the formula $\alpha_{j}\in\Lambda$ by the partial valuation $V[i]=v_{p}^{i}$. That is, $V[i][j]=v_{p}^{i}(\alpha_{j})$. This notation will be used in Algorithm 1, to be described in the sequel, also for subsets of $iPV(\Lambda)$.
Algorithm 1 (in Figure 4) for obtaining $iPLV(\Lambda)$.
Let $\Lambda\in FCS(\Omega)$, and consider the notation stipulated in Notation 9.6. Hence, $\Lambda=\{\alpha_{1},\ldots,\alpha_{n}\}$ and $iPV(\Lambda)=\{v_{p}^{1},\ldots,v_{p}^{m}\}$ (which can also be seen as sequences). Let $V:=iPLV(\Lambda)$. For each element $v_{p}^{i}$ of $V$, if $v_{p}^{i}(\alpha_{j})=\mathbf{U}$, determine if there is some $v_{p}^{i^{\prime}}$ such that $\mathsf{P}_{\Lambda}(v_{p}^{i},v_{p}^{i^{\prime}},\alpha_{j})$. This is represented by the function isSupported(V,i,j)(V,i,j) in Figure 4. If, for some $j$, isSupported(V,i,j)(V,i,j) is false (meaning no valuation in $V$ supports $v_{p}^{i}$ w.r.t. $\alpha_{j}$) then $v_{p}^{i}$ is removed from the candidate set for $iPLV(\Lambda)$ and will not be considered in subsequent algorithm iterations. After checking all the elements of $V$ using the elements of $V$ itself, let $IPV$ be the set of remaining valuations. If no valuation was removed, i.e., if $IPV=V$, the algorithm stops at this point and returns $IPV$ as being the set $iPLV(\Lambda)$ of partial level valuations. The set $IPV$ is the result of a refinement function applied to $V$, represented by refine(V)(V) in Figure 4. In case $IPV\neq V$, it means that $IPV$ is a proper subset of $V$ (i.e., some valuations were effectively removed in this round of the process). The set $IPV$ needs to be checked again with the function refine, to guarantee that every row is still supported. Then, the set $V$ is updated to be equal to $IPV$ (and so $V$ is, again, the universe of partial valuations to be checked), it is renumbered as a sequence $V[1],\ldots,V[m^{\prime}]$ (where $m^{\prime}<m$ is the cardinal or $IPV$) and then we repeat the process with the updated $V$ until no changes occur. The set $iPLV(\Lambda)$ is then defined to be the fixpoint $V=IPV$ (see Figure 4).
Figure 4: Algorithm for obtaining the set $iPLV(\Lambda)$.
Next result shows that, thanks to Algorithm 1, the set $iPLV(\Lambda)$ is well-defined, it is unique, and it is non-empty:
Proposition 9.7 (Existence and uniqueness of intuitionistic partial level valuations).
Let $\Lambda\in FCS(\Omega)$. Then, there exists a unique set $iPLV(\Lambda)\subseteq iPV(\Lambda)$ satisfying the condition of Remark 9.4:
$\begin{array}[]{lll}iPLV(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV(\Lambda)$}\big)\}.\end{array}$
Moreover, the set $iPLV(\Lambda)$ is non-empty, having at least $2^{k}$ elements, where $k$ is the number of different propositional variables occurring in $\Lambda$.
Proof.
Given $\Lambda\in FCS(\Omega)$, define the sequences $(\alpha_{1},\ldots,\alpha_{n})$ and $(v_{p}^{1},\ldots,v_{p}^{m})$ representing $\Lambda$ and $iPV(\Lambda)$ as in Notation 9.6, and then execute Algorithm 1 as described. Notice that $iPV(\Lambda)$ is a finite set, and the algorithm operates on a subset $V\subseteq iPV(\Lambda)$ in each round. If the algorithm does not stop in a given round, it deletes some elements from $V$ and restarts the process with a new subset that is strictly smaller than the previous one. This shows that the process will eventually terminate. There are two possible termination conditions: either every element of $V$ is eliminated, or a set $IPV=V$ is produced. In the second case, every element of $IPV$ is supported (as no removal is possible at this point); hence, $iPLV(\Lambda):=IPV$ satisfies the condition of Remark 9.4. In turn, the first possibility (arriving to a set $IPV=\emptyset$) is impossible by the following reason: every partial valuation $v:\Lambda\rightarrow V_{IPL}$ such that $v(\alpha)\in\{\mathbf{T},\mathbf{F}\}$ for every $\alpha\in\Lambda$ (i.e., every classical partial valuation) is never removed from $iPV(\Lambda)$ by means of Algorithm 1. Indeed, any classical valuation $v$ never satisfies that $v(\alpha_{j})=\mathbf{U}$ for some $j$, hence it will never be removed. Since there are exactly $2^{k}$ classical valuations in $iPV(\Lambda)$ (where $k$ is the number of different propositional variables occurring in $\Lambda$), the set $IPV$ produced by Algorithm 1 is non-empty, having at least $2^{k}$ elements. We have proven, therefore, the existence of a (non-empty) subset $IPV$ of $iPV(\Lambda)$ satisfying the condition of Remark 9.4.
To prove the uniqueness of such a set of partial valuations, observe that the order of the valuations $V[i]$ in the array $V$ is irrelevant for Algorithm 1. The algorithm’s execution does not depend on this initial ordering. In each round, every valuation $V[i]$ is tested against every formula $\alpha_{j}$ in $\Lambda$. If $V[i][j]=\mathbf{U}$, the function isSupported(V,i,j)(V,i,j) checks all other valuations $V[i^{\prime}]$ in $V$ to determine whether $\mathsf{P}_{\Lambda}(V[i],V[i^{\prime}],\alpha_{j})$ holds. If this condition fails for every $i^{\prime}$, $V[i]$ is removed in the next step. This procedure is clearly independent of the lexicographical order used in the formal presentation of Algorithm 1: that order was chosen solely to provide a precise description of the algorithm’s steps. In other words, regardless of the initial arrangement of $iPV(\Lambda)$, the first round will eliminate the same valuations and produce a unique set $V_{1}$ for the second stage. The same argument applies: any arrangement of $V_{1}$ yields the same subset $V_{2}$ after the second round. By induction, the algorithm’s output — the fixpoint $V_{s}=IPV$ reached in the final step $s$ — must be unique.
This completes the proof. ∎
Remark 9.8.
Observe that it is possible to have a set $iPLV(\Lambda)$ depending on $k$ different propositional variables $p_{1},\ldots,p_{k}$ having exactly $2^{k}$ elements. It is enough considering a formula $\alpha(p_{1},\ldots,p_{k})$ constructed by using exclusively conjunction and disjunction. Let $\Lambda$ be the set of subformulas of $\alpha$. Clearly, the truth-table $iPLV(\Lambda)$ for $\alpha$ only contains classical values $\mathbf{T}$ and $\mathbf{F}$, coinciding so with $iPV(\Lambda)$. It has, therefore, exactly $2^{k}$ elements. Moreover, by expanding this table with a column for $\alpha\to\alpha$, i.e., by considering $iPV(\Lambda\cup\{\alpha\to\alpha\})$, the last column only contains values $\mathbf{U}$ or $\mathbf{T}$. But then the rows assigning $\mathbf{U}$ to $\alpha\to\alpha$ are not supported by any other row, hence they are removed in the first round of Algorithm 1, arriving so to a set $iPLV(\Lambda\cup\{\alpha\to\alpha\})$ with $2^{k}$ elements.
@@
An intuitionistic truth-table in $\mathcal{M}_{IPL}$ is an exhaustive list of every and only intuitionistic partial level valuations for a given finite domain closed under subformulas.
@@
Example
In what follows, we present an example of a truth-table produced by Algorithm 1 which keep track of the dependencies of each partial valuation [^6] This example was automatically generated by a Rocq module for RNmatrices written by R. Leme, see https://github.com/renatoleme/Forest . The first table (Table 5(a)) is the full table for $\neg\neg(p\lor\neg p)$, a valid intuitionistic formula, and was obtained by applying the Nmatrix for $\mathbf{IPL}$ in the usual way.
(a) Initial table.
(b) First cycle of dependencies.
(c) Intermediate table.
(d) Second cycle of dependencies.
(e) Final table.
(f) Final dependency table.
Figure 5: Example of executions of refinement algorithm. Note that each partial valuation is identified by its ID, which is an index $n\in\mathbb{N}$. The dependency tables are obtained in two steps: first, find every $v$ such that there is some $\alpha$ such that $v(\alpha)=\mathbf{U}$. Then, for each of these partial valuations, find every valuation $w$ such that $w(\alpha)=\mathbf{F}$ and, for every $\beta$, $w(\beta)=\mathbf{T}$ whenever $v(\beta)=\mathbf{T}$.
In Figure 5, the tables on the left correspond to an execution of Algorithm 1 (see Figure 4), while the tables on the right represent the dependency tables. Whenever $v(\alpha)=\mathbf{U}$ occurs in a table on the left, the corresponding table on the right contains a set of supports (i.e., partial valuations $w$ such that $w(\alpha)=\mathbf{F}$ and $w(\beta)=\mathbf{T}$ whenever $v(\beta)=\mathbf{T}$) in the correspondent cell. If no such support exists for that case, then the corresponding partial valuation is removed in the subsequent execution of Algorithm 1, which can be observed in the transition between successive tables on the left.
By removing rows $3$, 44, and $6$ from Table 5(a), we obtain Table 5(b). In this new table, the row $1$, which was previously supported by the row $3$, no longer has that support available. Running the algorithm again with the updated table, we find that the row $1$ must now also be removed. This results in Table 5(e). At this point, Table 5(f) contains no occurrences of $\emptyset$, which means that every instance of $\mathbf{U}$ is now adequately supported. Hence, there are no more rows to remove and the algorithm is terminated. Finally, by inspecting the resulting Table 5(e), we can confirm that the formula is indeed valid.
Algorithm 1 is fundamentally the same as the one presented by Grätz ([12], Algorithm $3.4$), but is applied to the new RNmatrix for $\mathbf{IPL}$. As a result, it remains within the same complexity class using exponential space. However, it has a lower upper bound because the atomic valuation considers only two truth-values: $\mathbf{F}$ and $\mathbf{T}$. Section 10 provides a further discussion on this topic.
9.2 Soundness and completeness of the intuitionistic truth-tables
@@
Now, it will be shown that the generalized notion of truth-tables for IPL proposed in Subsection 9.1 is sound and complete. This will be established by first proving analyticity and co-analyticity with respect to level valuations in $\mathcal{M}_{IPL}$. To this end, and adapting [12, Definition 3.12], we introduce the intermediary notion of intuitionistic partial level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$.
@@
Let $\Lambda\in CS(\Omega)$ and $v_{p}\in iPV(\Lambda)$. Then, $\tilde{v}_{p}$ is an intuitionistic partial’ level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$ iff
@@
Definition 9.11 above is similar to Definition 9.3, except for two important aspects: (1) the domain $\Lambda$ of the partial valuation $\tilde{v}_{p}$ no longer needs to be finite, and (2) whenever there is a formula $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$, the support $w$ must be a level valuation.
This difference is central for proving both co-analyticity and analyticity. Aspect (1) is crucial for analyticity (Lemma 9.18), since we aim to prove that any domain closed under subformulas (possibly infinite) can be extended to a domain over the entire set of formulas. Aspect (2), in turn, is essential for co-analyticity (Lemma 9.13), as it relies on the previous result obtained in Proposition 8.21, which establishes the existence of a certain level valuation under appropriate circumstances.
The notion of $iPLV^{\prime}$ is well-suited for proof-theoretic purposes, but it is not sufficient for obtaining a decision procedure, since it depends on the set of level valuations. For this reason, we later prove that, for finite domains (which correspond to truth-tables), Definition 9.11 and Definition 9.3 are equivalent (see Proposition 9.19). This result, a direct consequence of Proposition 9.7, allows us to replace the notion of $iPLV^{\prime}$ with its finitary counterpart $iPLV$, showing that, in order to decide the validity of a formula (or argument), it is sufficient to restrict the analysis to the finite universe of its subformulas (see Theorem 9.22).
@@
Recall the predicate $\mathsf{P}_{\Lambda}(v,w,\alpha)$ introduced in Remark 9.4. Clearly,
$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big(\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}_{IPL}$}\big)\}.\end{array}$
@@
Let $v\in\mathcal{L}_{IPL}$ and $\tilde{v}_{p}:=v_{|\Lambda}$. Clearly, $\tilde{v}_{p}\in iPV(\Lambda)$. Let $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$. Hence, $v(\alpha)=\mathbf{U}$. By Proposition 8.21, there exists $w\in\mathcal{L}_{IPL}$ such that $w(\alpha)=\mathbf{F}$ and $w(\beta)=\mathbf{T}$ for every $\beta$ such that $v(\beta)=\mathbf{T}$. In particular, $\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)$ holds. From this, $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$ by Remark 9.12. ∎
A key concern regarding truth-tables is whether a given partial valuation can always be extended to a full valuation, i.e., whether a domain closed under subformulas can be extended to a complete domain. This property, called analyticity, is central for proving the completeness of the method (see Theorem 9.22). To establish analyticity, we start with a domain over a set closed under subformulas and conservatively add new formulas to this domain. At the limit, we obtain the function given by the union of all such extensions, which corresponds to a function defined over the set of all formulas. However, a formula can be evaluated only if all of its subformulas are already evaluated. To ensure that this condition always holds, we consider a suitable ordering on the set of formulas. The technique will be entirely analogous to the case of S4, recall Remark 6.15.
Definition 9.14.
@@
Given the ordinal $\omega$, let $\omega^{2}:=\omega\cdot\omega$ be its ordinal multiplication. Observe that it is possible to define an enumeration $\alpha_{1},\alpha_{2},\ldots\alpha_{m},\ldots$ (for $m\in\omega^{2}$) of $For(\Omega)$ such that $\textsf{c}(\alpha_{i})\leq\textsf{c}(\alpha_{j})$ if $i\leq j$ and, for every $i$ such that $\textsf{c}(\alpha_{i})>0$, if $\beta$ is a strict subformula of $\alpha_{i}$ then $\beta=\alpha_{j}$ for some $j<i$. Such an enumeration of $For(\Omega)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 0 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
Based on the enumeration of $For(\Omega)$ given in the previous remark, consider the following sets:
Definition 9.16.
Let $\alpha_{1},\alpha_{2},\ldots\alpha_{m},\ldots$ (for $m\in\omega^{2}$) be an enumeration of $For(\Omega)$ as in Remark 9.15. For every $n,m\in\omega$ let $\bar{\Lambda}_{n}:=\Lambda\cup For(\Omega)_{n}$; $\bar{\Lambda}_{n}^{0}:=\bar{\Lambda}_{n}$; and $\bar{\Lambda}_{n}^{m+1}=\bar{\Lambda}_{n}^{m}\cup\{\alpha_{\omega\cdot(n+1)+m}\}$.
Remark 9.17.
By definition, $\bar{\Lambda}_{n}^{m+1}=\bar{\Lambda}_{n}\cup\{\alpha_{\omega\cdot(n+1)},\alpha_{\omega\cdot(n+1)+1},\ldots,\alpha_{\omega\cdot(n+1)+m}\}$, and $\alpha_{\omega\cdot(n+1)+m}\in\Lambda$ if and only if $\bar{\Lambda}_{n}^{m+1}=\bar{\Lambda}_{n}^{m}$. Hence, $\bar{\Lambda}_{n}^{m+1}$ is obtained from $\bar{\Lambda}_{n}$ by adding the first $m+1$ formulas (of the given enumeration) with complexity $n+1$, and so $\bar{\Lambda}_{n}^{m}$ adds at most $m$ formulas to $\bar{\Lambda}_{n}$, for $m\in\omega$. It is worth noting that $\bar{\Lambda}_{n}^{m}\in CS(\Omega)$ for every $n,m\in\omega$; in particular, $\bar{\Lambda}_{n}\in CS(\Omega)$ for every $n\in\omega$. Clearly $\bar{\Lambda}_{n}^{m}\subseteq\bar{\Lambda}_{n}^{m+1}$, $\bar{\Lambda}_{n+1}=\bigcup_{m\in\omega}\bar{\Lambda}_{n}^{m}$, and $For(\Omega)=\bigcup_{n\in\omega}\bar{\Lambda}_{n}$.
@@
It is an adaptation of the proof of the Analyticity lemma for $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ (recall Lemma 6.18) to $\mathcal{R}(\mathcal{M}_{IPL})$.